OpenStudy (anonymous):

The sum of the real roots of the equation Ix-2I^2 + Ix-2I-2=0 is - a)2 c)4 b)6 d) 8

OpenStudy (anonymous):

@nitz

OpenStudy (anonymous):

\[\left| x-2 \right|^{2}+\left| x-2 \right|-2\]

OpenStudy (anonymous):

=0

OpenStudy (anonymous):

x=3,1,0,4

OpenStudy (anonymous):

is it?

OpenStudy (anonymous):

ans. 4 (sum of real roots)

OpenStudy (anonymous):

see what i think i can explain you

OpenStudy (anonymous):

put \[\left| x-2 \right|=t\]

OpenStudy (anonymous):

the equation becomes; \[t ^{2}+t-2=0\]

OpenStudy (anonymous):

t=1 and t=-2

OpenStudy (anonymous):

got it.

OpenStudy (anonymous):

when t=1; \[\left| x-2 \right|=1\]

OpenStudy (anonymous):

ok ok i got it.

OpenStudy (anonymous):

x-2=1 and x-2=-1 x=3 and x=1 so the sum is 4 in second case: when t=-2 \[\left| x-2 \right|=-2\]

OpenStudy (anonymous):

x-2=-2 and x-2=2 x=0 and x=4 sum is 4 again

OpenStudy (anonymous):

nitz wait

OpenStudy (anonymous):

mod of x-2 cant be equal to a negative no. So it will be rejected??

OpenStudy (anonymous):

ya right it will be rejected as mod(X) is always positive good haan

OpenStudy (anonymous):

ok thanks need you on some more.

OpenStudy (anonymous):

ya sure